OpenStudy (anonymous):

Hi, my question is how do you square a quantity like (2sqrt11)^2

OpenStudy (johnweldon1993):

\[\large (2\sqrt{11})^2\] we can write it as \[\large 2\sqrt{11} \times 2\sqrt{11}\] So what is 2 times 2? and what is √11 times √11?

OpenStudy (anonymous):

2x2=4, I'm not sure what the square root of 11 x square root of 11 is

OpenStudy (johnweldon1993):

Well...we know √11 times √11 can be written as \[\large \sqrt{11}^2\] right? "The square of a square root" they cancel each other out So the square root on 11 times the square root of 11 = 11 so this question becomes \(\large 4\times 11\)

OpenStudy (anonymous):

but wouldn't it go like 4+2sqrt11+2sqrt11+sqrt11^2?

OpenStudy (johnweldon1993):

Nope....if it were \[\large (2 + \sqrt{11})^2\] it would be that case! because you would distribute it out! however this is just multiplication \[\large (2\sqrt{11})^2\]

OpenStudy (anonymous):

oh ok, that makes sense.

OpenStudy (anonymous):

Another question I have is that it says to multiply and simplify. (7 3sqrt4)(6 3sqrt2)